The solubility of AgCN (s) in 0.200M NH3 (aq) is 8.8×10−6mol/L.

Calculate Ksp for AgCN.

ANSWER Ksp = 1.2×10−16
Answer #1

When we dissolve AgCn in NH3, it forms a complex as [Ag(NH3)2]+

Given that solubility of AgCN in NH3 is 8.8 *10-6 mol/L

and reaction is:

AgCN(s) + 2NH3(aq) <=> [Ag(NH3)2]+(aq) + CN-(aq)

Solubility =  [Ag(NH3)2]+ = [CN-] = 8.8 *10-6

We have

AgCN(s) <=> Ag+(aq) + CN-(aq) .... Ksp =?

Ag+(aq) +2NH3(aq)---> [Ag(NH3)2]+(aq)...Kf = 1.6 *107 ... From Literature

Overall reaction

AgCN(s) + 2NH3(aq) <=> [Ag(NH3)2]+(aq) + CN-(aq) .....K = Ksp*Kf

K =  [Ag(NH3)2]+ [CN-] / [NH3]2

Solids have activity =1. So, not included in reaction

K = (8.8 *10-6)2 / (0.200) = 1.9 *10-9

Ksp = K/Kf = (1.9 *10-9) / ( 1.6 *107) = 1.2 *10-16